Home‎ > ‎Reports‎ > ‎Crew099‎ > ‎

day10


Crew 99 Log Book for February 7, 2011

Photos of the day





Commander's Check-In Report

Donna Viola Reporting

Crew Physical Status: good


Time Departed/Returned from EVA: EVA #14, 11:15-12:30; EVA #15,

14:15-14:45; EVA #16, 15:00-17:00


Brief Narrative of Field Mission Results: Endolith sample collection;

EVA planner; Collecting more ATV waypoints


EVA Data/Interpretations: as per EVA reports


Engineering/Hab Maintenance: Nothing major to report.


Report Transmission Schedule: Engineering, EVA #14, EVA #15, EVA #16,

Commander's, Cook's. By 8:30pm


Plans for Tomorrow: to be decided


Inventory: none


Support Requested: Would it be permissible for us to make a trip to Grand Junction with the Hab car on Friday? We have helium tanks for the balloon launch project that need to be returned there, and we are concerned about those fitting in the rental car with us on Saturday.


Miscellaneous: N/A



Commander's Report

Donna Viola Reporting

Exhausted from a tiring game of football yesterday, the team slept well last night - and awoke for a busy day of EVAs!

In the morning, Sukrit and Max joined me to go hunting for endoliths north of the Hab.

Both took turns with the rock hammer, and helped me find some really good specimens. In the afternoon, Andie and Kevin went out to test the

EVA planner again. After they returned, Heidi, Kevin, and I went to track down some more coordinates for the ATV roads - and saw some pretty stunning vistas.

We're hoping to compile a database containing those points to help future crews navigate the area.


Engineering Report

Kevin Newman Reporting


Generator/Electricity:


· Kitty (Diesel Generator) (Used? Yes/No): Yes


· Honey (Gas Generator) (Used? Yes/No): No


· Battery Charge Level: Checked by DG


o 10.59 V Everstart


o 12.37 V Armor plate


o 13.32 V Green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run


Fuel Status:


· Diesel (Full/0.75/0.5/0.25/Empty): 0.25


· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.74


· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 17 Gallons


Notes/Comments:


Oil:


· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG


· Honey Oil Quantity (# of Quarts in storage): Leave it to DG


· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench


· Car Oil Quantity (# of Quarts in storage):


Water Status:


· Outside Potable Water Tank Level (inches from bottom): 17


· Trailer Potable Water Tank Level (inches from bottom): 5/6 Full


· Hab Potable Water Tank Level (inches from bottom): 12


· Potable Water Meter Reading: 417928


Notes/Comments:


GreenHab:


· Condition of plants in Tank 1: Nothing


· Condition of plants in Tank 2: Ok


· Condition of plants in Tank 3: Good


· Greenhab notes/ comments:


Observatory:


· Telescope Used? (Yes/no) No


Observatory Notes:


Transportation:


Hab Car (NOT RENTAL)


· Used (yes/no): No


· Oil Checked (yes/no): No


· Tire Status: Not checked


· Odometer at end of day: 123611.6


· Notes/Comments on Hab Car:


ATV 1 (Opportunity):


· Used: Yes


· Oil Checked: No


· Fuel Consumed: Replaced


· Tire Status: Filled


· Comments:


ATV 2 (Spirit):


· Used: Yes


· Oil Checked: No


· Fuel Consumed: Replaced


· Tire Status: Filled


· Comments:


ATV 3 (Viking I):


· Used: No


· Oil Checked: No


· Fuel Consumed: None


· Tire Status: Not Checked


· Comments:


ATV 4 (Viking II):


· Used: Yes


· Oil Checked: No


· Fuel Consumed: Replaced


· Tire Status: Filled


· Comments:


Heating and Ventilation:


· Thermostat setting upstairs (Farenheit): 66


· Actual temperature upstairs (Farenheit): 80


· Thermostat setting downstairs (Farenheit): 65


· Actual temperature downstairs (Farenheit): 72


Computers/Networking Infrastructure:


Notes/comments on internet/computers?: Behaving well

General Engineering

1. Notes/Comments:Someone asked for pictures of the ATV covers, so they are attached.
Clear Skies!


EVA 14 Report

Max Fagin Reporting

Time: 11:00-12:30


Crew Members: Donna Viola, Max Fagin, Sukrit Ranjan


Site Location:  N 38° 26' 32.9"   :    W 110° 47' 44.1"


Transit Mode: ATV+Foot


Objectives: Endolith sample collection.


Results: A rock outcropping by a dried stream bed yielded several endolith samples.  Nearby gypsum patches on the ground were also found, indicating that the region had been submerged in water in the past (see photos).


Lessons Learned: It took us 10 minutes to walk across the lake bed, select a rock, break the rock open, identify the endoliths, select an appropriate sample, photograph it, bag it, and move on to the next rock.  As wonderful as Spirit and Opportunity are, it probably would have taken them a month to do the same tasks.  A human with a $200 EVA suit and a $12 rock hammer is 4000 times more efficient than a rover with a multi-million dollar instrument package.  We've just got to get us there first.


Dry lake beds are a great place to look for life, and gypsum is a great way to identify dry lake beds.  But this was only clear after Donna had explained it to the rest of us.  The first humans on Mars should be trained to recognize the geologic signs of water as quickly as they do their mother's face.


EVA 15 Report

Andie Gilkey Reporting

Time: 14:15-14:45


Crew Members Andie Gilkey, Kevin Newman


Site Location:  Thielges Flats

Transit Mode: Foot


Objectives: Use EVA mission planner to plan mission and contingency while in the field


Results: Mission planner was successfully used to find the site and re-plan in the field


Lessons Learned:  It is difficult to see the iPad screen in the bright sunlight


EVA 16 Report

Heidi Beemer Reporting

Time: 1500- 1700

Crew Members Heidi Beemer, Donna Viola, Kevin Newman

Site Location:  Lith Canyon

 Transit Mode: ATV

Objectives: Part: Find Ptolemy Road and explore Lith Canyon on foot

Results: Ptolemy Road could not be found. We are still disconnected from the North Eastern part of the map. Walkway into Lith Canyon was found and we explored around. No Dinosaurs.

Lessons Learned:  After searching for a road that does not exist you start imagining roads that are not really there.


Journalist Report

Sukrit Sanjan Reporting

We woke late today after a rambunctious day of football. In the morning, Donna, Max and Sukrit went on EVA to collect endoliths. We identified thriving colonies. It was amazing to see how life can thrive in harsh environments. We also saw gypsum, a signature of past water flow.

After, Andie and Kevin went on EVA planner EVA, followed by another mapping EVA in conjunction with Heidi. For dinner, we had an extremely filling dinner of mushroom pilaf, and are now settling in for the evening.


Cook's Report

Heidi Beemer Reporting

1a) Today was

-A non-cooking day

1b) Was today a special day celebrated at the main meal with special

food or activity?

Wave all of your fingers at your neighbor day.

2)  List the foods served at today's main meal (usually the evening

meal), giving full names of each.

Gourmet Reserves Mushroom Pilaf

3) List any main meal foods not finished by the crew (leftovers stored

for later use or discarded)  (Example:  2 servings potato soup stored;

2 servings rice burned/inedible, discarded)

- We had two servings of leftover.

4)  Recipes for anything more complicated than following package

directions exactly, or rehydrating.

-None

5)  Comments/questions for Kim and Jean, the food study investigators

- None